OpenStudy (anonymous):

Find the distance between the points (1, 3) and (-7, 2).

OpenStudy (anonymous):

ANSWERS: \[\sqrt{65}\] \[\sqrt{37}\] \[\sqrt{16}\]

OpenStudy (anonymous):

@Arnab09

OpenStudy (anonymous):

distance between 2 points= sqrt((x1-x2)^2 + (y1-y2) ^2) so in this case, sqrt [(1+7)^2+(3-2)^2]=sqrt65

OpenStudy (anonymous):

thanks!